OpenStudy (anonymous):

How can you solve the equation y - 2.9 = 5.3 ? choices: Subtract 5.3 from both sides. Subtract 2.9 from both sides. Add 2.9 to both sides. Add 5.3 to both sides.

OpenStudy (anonymous):

C) add 2.9 to both sides

OpenStudy (anonymous):

You have to cancel the " -2.9" on the Left side. So add "2.9" to it -2.9 +2.9 = 0 However this is an equation, so what you do one one side, you have to repeat the same on the other side as well. So add 2.9 to the 5.3 on the right side also so, y -2.9 (+ 2.9)= 5.3 (+ 2.9) y = 5.3 +2.9 y = 8.2 Answer : "Add 2.9 to both sides."